Ellen wishes to mix candy worth $1.69 per pound with candy worth $3.91 per pound to for 30 pounds of a mixture worth $2.73 per pound. How many pounds of the more expensive candy should she use? Round your answers to the nearest pound.
Answer by richwmiller(17219)

test as a category is for testing how the website works.
This problem belongs in word problems mixtures
1.69x+3.91y=30*(2.73)
x+y=30
x=15.9459, y=14.0541
14 pounds of 3.91 candy